# cmaes : A Simple yet Practical Python Library for CMA-ES

### v1

@@ Abstract.
$*{}^{*}$ These authors contributed equally. This work was conducted while Masashi Shibata was affiliated with CyberAgent.
@@ 2. CMA-ES
(7) $\sigma^{(g+1)}=\sigma^{(g)}\exp\left(1,\frac{c_{\sigma}}{d_{\sigma}}\left(\frac{\|p_{\sigma}^{(g+1)}\|}{\mathbb{E}[\|\mathcal{N}(0,I)\|]}-1\right)\right),$
@@
(8) $\ +c_{1}\underbrace{\left[p_{c}^{(g+1)}\left(p_{c}^{(g+1)}\right)^{\top}-C^{(g)}\right]}_{\text{rank-one update}}+c_{\mu}\underbrace{\sum_{i=1}^{\lambda}w_{i}^{\circ}\left[y_{i:\lambda}y_{i:\lambda}^{\top}-C^{(g)}\right]}_{\text{rank-$\mu$ update}},$
@@ 3.2. Getting Started
To illustrate its basic functionality, Listing LABEL:list:ask-and-tell-api shows a simple example using cmaes. The library employs an ask-and-tell interface (Collette et al., 2013). After initializing CMA(normal-$\cdot$) with initial mean and sigma, the process involves the following steps: (1) Candidate solutions are generated using the ask() method. (2) The sampled solutions are then evaluated by an objective function. (3) The distribution parameters are updated based on the solutions and their objective function values using the tell(normal-$\cdot$) method.
@@
This ask-and-tell interface, by decoupling the optimizer from the objective function, adds flexibility to the library. For example, it is easy to implement CMA-ES with restart strategies, such as IPOP-CMA-ES (Auger and Hansen, 2005) or BIPOP-CMA-ES (Hansen, 2009). Listing LABEL:list:ipop-cma-es provides an example code for IPOP-CMA-ES. In this example, the algorithm restarts with a new mean vector and a doubled population size when the stopping conditions are met (i.e., when should_stop() returns True). These stopping conditions are based on those used in pycma (Hansen et al., 2019). Additionally, we can specify box constraints by adding the argument, for example, bounds=np.array([[-3,3],[-3,3]]) to CMA(normal-$\cdot$); if a solution is not feasible, a resampling method is used, which is reasonable when the optimum solution is not very close to the constraint.
@@ 4.1. Multimodal and Noisy Problems
Users can run LRA-CMA simply by adding the argument lr_adapt=True during the initialization of CMA(normal-$\cdot$).
@@ 4.2. Transfer Learning
Listing LABEL:list:ws-cma shows an example code using WS-CMA. We first need to prepare an array (source_solutions in the code) whose element is a tuple $(x,f(x))$. By using the method get_warm_start_mgd(normal-$\cdot$) with the created source_solutions, we can obtain the promising distribution parameters of the mean vector, the step-size, and the covariance matrix. WS-CMA is then performed by initializing CMA(normal-$\cdot$) with these parameters.

### v2

@@ Abstract.
These authors contributed equally. This work was conducted while Masashi Shibata was affiliated with CyberAgent.
@@ 2. CMA-ES
(7) $\sigma^{(g+1)}=\sigma^{(g)}\exp\left(\frac{c_{\sigma}}{d_{\sigma}}\left(\frac{\|p_{\sigma}^{(g+1)}\|}{\mathbb{E}[\|\mathcal{N}(0,I)\|]}-1\right)\right),$
@@
(8) $\ +c_{1}\underbrace{\left[p_{c}^{(g+1)}\left(p_{c}^{(g+1)}\right)^{\top}-C^{(g)}\right]}_{\text{rank-one update}}+c_{\mu}\underbrace{\left[\sum_{i=1}^{\lambda}\left(w_{i}^{\circ}y_{i:\lambda}y_{i:\lambda}^{\top}-w_{i}C^{(g)}\right)\right]}_{\text{rank-$\mu$ update}},$
@@ 3.2. Getting Started
To illustrate its basic functionality, Listing LABEL:list:ask-and-tell-api shows a simple example using cmaes. The library employs an ask-and-tell interface (Collette et al., 2013). After initializing CMA($\cdot$) with initial mean and sigma, the process involves the following steps: (1) Candidate solutions are generated using the ask() method. (2) The sampled solutions are then evaluated by an objective function. (3) The distribution parameters are updated based on the solutions and their objective function values using the tell($\cdot$) method.
@@
This ask-and-tell interface, by decoupling the optimizer from the objective function, adds flexibility to the library. For example, it is easy to implement CMA-ES with restart strategies, such as IPOP-CMA-ES (Auger and Hansen, 2005) or BIPOP-CMA-ES (Hansen, 2009). Listing LABEL:list:ipop-cma-es provides an example code for IPOP-CMA-ES. In this example, the algorithm restarts with a new mean vector and a doubled population size when the stopping conditions are met (i.e., when should_stop() returns True). These stopping conditions are based on those used in pycma (Hansen et al., 2019). Additionally, we can specify box constraints by adding the argument, for example, bounds=np.array([[-3,3],[-3,3]]) to CMA($\cdot$); if a solution is not feasible, a resampling method is used, which is reasonable when the optimum solution is not very close to the constraint.
@@ 4.1. Multimodal and Noisy Problems
Users can run LRA-CMA simply by adding the argument lr_adapt=True during the initialization of CMA($\cdot$).
@@ 4.2. Transfer Learning
Listing LABEL:list:ws-cma shows an example code using WS-CMA. We first need to prepare an array (source_solutions in the code) whose element is a tuple $(x,f(x))$. By using the method get_warm_start_mgd($\cdot$) with the created source_solutions, we can obtain the promising distribution parameters of the mean vector, the step-size, and the covariance matrix. WS-CMA is then performed by initializing CMA($\cdot$) with these parameters.